The production function of given firm Is given by Q =200sqrtLK  and the.

associated cost function is given by 2L + 5K = 1000

A. Calculate the level of labor and capital input at the optimum level of production

B. Calculate the MRTSL,and MRTSkL

c. What is the level of output at the least cost combination of labor and capital inputs?

1
Expert's answer
2021-09-30T08:17:13-0400

A. The condition of optimum for producer is MP(L)/w=MP(K)/r (****); MP(L)=Q'(L)=1/2×200×L^(-1/2)×K^1/2; MP(K)=Q'(K)=1/2×200×L^1/2×K^(-1/2). w and r are prices for L and K, so they can be found from equation of costs: w= 2; r=5. Then, after simplification K=2/5 L. Applying this in the cost equation, 2L+5×2/5L=1000; L=250, K=100.

B)MRTS L,K =MP(L)/MP(K); Sipmlifying the relevant expressions from (A), we get

MRTS L K=K/L, MRTS K, L=L/K.

C) Q=200×sqrt(250×100)=31623